Circular No. 5875

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SGR 1806–20 C. Kouveliotou, J. M. Horack, G. J. Fishman, C. A. Meegan, and W. S.

Paciesas report for the BATSE Team: “We have detected repeated soft γray emission from a region centered at α = 18^{h}16^{m}, δ = −20^{o} (equinox 2000.0) with an average error radius of ∼ 6^{o}. Three triggers were recorded on Sept. 29.40280, 29.85163, and 29.96912 UT. All triggers are similar; each consists of a single pulse of duration ∼ 75, 200, and 40 ms at full width zero amplitude, respectively. Their spectra are consistent with a typical Soft Gamma Repeater (SGR) spectrum; no emission is detected above 100 keV. Our results are consistent with the location of the SGR 1806–20 (Atteia et al. 1987, Ap.J. 320, L105). The high frequency of the triggers (three in ∼ 14 hr) is consistent with the emission pattern of this source, which is known to emit multiple episodes per day (Laros et al. 1987, Ap.J. 320, L111). Our results indicate that SGR 1806–20 has become active again. Follow-up observations of the source region at other wavelengths are strongly encouraged.”

PLUTO A. Gilmore and P. Kilmartin, Mount John University Observatory,

report that observations of star P20, Pluto, and Charon, were made during Oct. 3.358-3.386 UT with the 0.6-m f/13 Cassegrain reflector (+ single channel photometer No. 1 + Cousins R filter + 30′′ aperture), revealing no occultation by Charon (cf. IAUC 5872). Every 2-3 min, integrations were interrupted for 5-10 s to check tracking. There was thin cirrus near the region and possibly over it. The net count rate was ∼ 100 Hz at the start and declined to ∼ 50 Hz at the end of observation; no changes occurred in count rate other than those attributable to the increasing air mass (to > 6 air masses). An integration time of 0.5 s was used because of the low rate.

NOVA SAGITTARII 1993 Visual magnitude estimates (cf. IAUC 5868): Sept. 23.12 UT, 9.1

(R. Royer, Lakewood, CA); 23.72, 9.1 (D. Overbeek, Edenvale, S. Africa); 25.83, 8.7 (A. Pereira, Cabo da Roca, Portugal); 26.41, 9.5 (A. F. Jones, Nelson, New Zealand); 26.83, 9.4 (Pereira); 27.41, 9.1 (Jones); 27.82, 9.6 (Pereira); 28.37, 9.2 (Jones); Oct. 3.83, 10.3 (Pereira).

1993 October 5 Daniel W. E. Green